A rectangular pool is surrounded by a walk 4 meters wide. The pool is 6 meters longer than its width. If the total area is 576 square meters more than the area of the pool, find the dimensions of the pool.
Answer by ankor@dixie-net.com(22740)

A rectangular pool is surrounded by a walk 4 meters wide. The pool is 6 meters longer than its width. If the total area is 576 square meters more than the area of the pool, find the dimensions of the pool.
:
Let the width of the pool = x
Then the length will = (x+6)
:
Draw a diagram of this with the 4'walkway. It will be apparent that the over all
dimensions will be (x+8) by (x+6+8) or (x+14)
:
Total area - pool area = 576
(x+14)*(x+8) - x(x+6) = 576
:
x^2 + 22x + 112 - (x^2 + 6x) = 576
:
x^2 + 22x + 112 - x^2 - 6x - 576 = 0
:
22x - 6x + 112 - 576 = 0
:
16x - 464 = 0
:
16x = 464
:
x = 464/16
:
x = 29 ft is the width of the pool, length: 29+6 = 35 ft
:
:
Check our solution of x=29
Overall - pool = 576
(37*43) - (29*35) =
1591 - 1015 = 576